Hi could you please explain how to use substitution through this problem? I have tried numerous times and i am not quite sure if i have done it correctly.
x-2y=1
6x-5y=20

Answer by rfer(16322)

x=2y+1
6(2y+1)-5y=20
12y+6-5y=20
7y=14
y=2
------------------
x-2(2)=1
x=5
------------------
6*5-5*2=20